I have two rectangles that I need help find if they are similar and if they are find there scale factor.One is 14cm by 6cm by 8cm and the other one is 7cm by 3cm by 4cm. How do I do this, Can you help me please?
Answer by mananth(16946)

Find the ratio of their corresponding sides.
if the ratios are equal then the rectangles are similar
14,6,8
7,3,4
14/7 = 6/3 = 8/4
hence similar